OpenStudy (anonymous):

graph tan x/3...I know the process. I need some guidance. :)

OpenStudy (anonymous):

What guidance? Where are you stuck?

OpenStudy (anonymous):

I am not sure how to graph with my calculator. I can make a list to sketch, but not sure I am doing it correctly.

OpenStudy (anonymous):

Have you studied calculus? Do you know how to use derivatives to find local extrema and inflection points? Can you find the zeros of that function? Or are you supposed to graph this with a graphing calculator?

OpenStudy (anonymous):

No I am in non-cac. based Trig for Physics.

OpenStudy (anonymous):

I used a calculator. I saw sin as 0. I thought.

OpenStudy (anonymous):

Yes I can find the zeros of this function.

OpenStudy (anonymous):

I have a trig. table chart.

OpenStudy (anonymous):

Now, let me help you. If you had to graph tan(x), you know that it has two asymptotes, that is where tan(x) becomes undefined. And that is at \[-\frac{ \pi }{ 2 }\]and\[\frac{ \pi }{ 2 }\]But since you have x/3, you divide the above values by three and you get the numbers where tan(x/3) becomes undefined. You plot em and you start graphing like this: